# Hierarchical Clustering for Conditional Diffusion in Image Generation

### v1

@@ Appendix A Detailed Formulation of TreeVAE
The TreeVAE is a novel deep generative model proposed by Manduchi et al. (2023). Built upon the VAE framework originally introduced by Kingma & Welling (2014), TreeVAEs inherently possess stochastic latent variables. However, unlike traditional VAEs, TreeVAEs organize these latent variables in a learnable binary tree structure , enabling them to capture complex hierarchical relationships.
@@
In the TreeVAE framework, the tree hierarchically divides the data, yielding separate stochastic embeddings for each node. This division into nodes induces a probability distribution, allowing each sample to navigate through the nodes of the tree in a probabilistic manner, from the root to the leaf. Ideally, high-variance features should be partitioned at earlier stages in the tree, while deeper nodes encapsulate more detailed concepts. The flexible tree structure is learned during training and is specific to the data distribution. This structure allows the incorporation of sample-specific probability distributions over the different paths in the tree, as further explained in Section A.1. In this manner, TreeVAEs contribute to the generation of comprehensive hierarchical data representations.
@@
TreeVAE combines elements from both LadderVAE (Sønderby et al., 2016) and ANT (Tanno et al., 2019), aiming to create a VAE-based model capable of performing hierarchical clustering via the latent variables. Similar to LadderVAE, the inference and the generative model share the same top-down hierarchical structure. In fact, if we isolate a path from the root to any leaf in the tree structure of TreeVAE, we obtain an instance of the top-down model seen in LadderVAE. On the other side, both Adaptive Neural Trees and TreeVAE engage in representation and architecture learning during training. However, while ANTs are tailored for regression and classification tasks, TreeVAE’s focus lies in hierarchical clustering and generative modeling. This allows TreeVAEs to generate new class-specific data samples based on the latent embeddings of the leaves from the tree.
A.1 Model Formulation
@@
TreeVAE comprises an inference model and a generative model, which share the global structure of the binary tree $\mathcal{T}$. Following Manduchi et al. (2023), the tree structure is learned during training using dataset $\bm{X}$ and a predetermined maximum tree depth denoted as $H$. Specifically, the tree structure entails the set of nodes $\mathbb{V}=\{0,...,V\}$, the subset of leaves $\mathbb{L}\subset\mathbb{V}$, and the set of edges $\mathcal{E}$.
@@
The transformations and routers are learned during the model training process, as further elaborated in Section A.2, which comprehensively covers all aspects of model training. The exact parametrizations of the transformations and routers vary depending on whether they are applied in the context of the inference or generative model. Subsequently, the specifics of both of these models will be explored.
@@
Following the generation of latent representations for each node $i$ in the tree, the decision regarding traversal to the left or right child node are determined based on the sampled $\bm{z}_{i}$ value. These decisions, denoted by $\mathrm{c}_{i}$, for each non-leaf node $i$ are guided by routers, which are neural network functions linked to the generative model and defined as $\left\{r_{p,i}\mid i\in\mathbb{V}\backslash\mathbb{L}\right\}$. Here, $\mathrm{c}_{i}=0$ indicates the selection of the left child of the internal node $i$, aligning with the design proposed by Manduchi et al. (2023). As mentioned before, these decisions follow a Bernoulli distribution, $\mathrm{c}_{i}\mid\bm{z}_{i}\sim\operatorname{Ber}\left(r_{p,i}\left(\bm{z}_{i}\right)\right),$ which, in turn, induces the following path probability between a node $i$ and its parent node $pa(i)$:
@@
To conclude the generative process, $\bm{x}$ is obtained based on the latent embeddings of a selected leaf $l$. Nevertheless, assumptions on the distribution of the inputs are required. For real-valued $\bm{x}$, such as in colored datasets, Manduchi et al. (2023) assume a Gaussian distribution. For grayscale images, they consider the Bernoulli distribution. Thus,
@@
where $\left\{\mu_{x,l},\sigma_{x,l}\mid l\in\mathbb{L}\right\}$ or $\left\{\mu_{x,l}\mid l\in\mathbb{L}\right\}$, respectively, are implemented as leaf-specific neural networks called “decoders”. Typically, a simplification is made by assuming $\sigma_{x,l}^{2}\left(\bm{z}_{l}\right)=\bm{I}$ for convenience.
@@
The inference model of TreeVAE, depicted on the left side of Figure 7, introduces a deterministic bottom-up pass to incorporate the conditioning on $\bm{x}$, distinguishing it from the generative model. This bottom-up process is akin to the framework proposed by Sønderby et al. (2016). In this hierarchical structure, the bottom-up deterministic variables depend on each other via a series of neural network operations, specifically MLP of the same architecture as the transformation MLP defined previously,
@@
The first deterministic variable in this chain, denoted as $\mathbf{d}_{H}$, serves as the output of an encoder neural network. This encoder accepts the input image $\bm{x}$ and transforms it into the flattened, low-dimensional vector $\mathbf{d}_{H}$. $H$ corresponds to both the number of deterministic variables involved in the bottom-up process and the maximum depth to which the tree structure can grow during training as further explained in Section A.2.
The tree structure is shared between the inference and the generative model, though adjustments are made to the node-specific parameterizations of the Gaussian distributions. Notably, in the inference phase, information is introduced through conditioning on $\bm{x}$, influencing the distribution of latent embeddings across all nodes. Hence, similar to Sønderby et al. (2016), the means $\bm{\mu}_{q,i}$ and variances $\bm{\sigma}_{q,i}^{2}$ of the variational posterior distribution for each node $i$ are calculated using dense linear network layers conditioned on the deterministic variable of the same depth as the node $i$:
@@
Finally, given the variational distributions of the latent embeddings equation 18 & equation 19 and the variational distributions of the decisions equation 20, Manduchi et al. (2023) construct the variational posterior distribution of the latent embeddings and paths:
$q\left(\bm{z}_{\mathcal{P}_{l}},\mathcal{P}_{l}\mid\bm{x}\right)=q\left(\bm{z}_{0}\mid\bm{x}\right)\prod_{i\in\mathcal{P}_{l}\backslash\{0\}}q\left(\mathrm{c}_{pa(i)\rightarrow i}\mid\bm{x}\right)q\left(\bm{z}_{i}\mid\bm{z}_{pa(i)}\right).$ (21)
A.2 Model Training
TreeVAE entails the training of various components, involving both the parameters of the neural network layers present in the inference and generative models, as detailed in Section A.1, and the binary tree structure $\mathcal{T}$. The training process alternates between optimizing the model parameters with a fixed tree structure and expanding the tree.
During a training iteration given the current tree structure, the neural layer parameters are optimized. These include the parameters of the inference model and the generative model, encompassing the encoder ($\mu_{q,0},\sigma_{q,0}$), the bottom-up MLPs from equation 14, the dense linear layers from equation 15 & equation 16, the transformations ($\left(\left\{\left(\mu_{p,i},\sigma_{p,i}\right),\left(\mu_{q,i},\sigma_{q,i}\right)\mid i\in\mathbb{V}\backslash\{0\}\right\}\right)$, the routers ($\left(\left\{r_{p,i},r_{q,i}\mid i\in\mathbb{V}\backslash\mathbb{L}\right\}\right)$, and the decoders ($\left(\left\{\mu_{x,l},\sigma_{x,l}\mid l\in\mathbb{L}\right\}\right)$. The objective in training these parameters is to maximize the likelihood of the data given the learned tree structure $\mathcal{T}$, denoted as $p(\bm{x}\mid\mathcal{T})$. This corresponds to modeling the distribution of real data via the hierarchical latent embeddings. To compute $p(\bm{x}\mid\mathcal{T})$, the latent embeddings must be marginalized out from the joint distribution:
@@ A.2 Model Training
Similar to other Variational Autoencoders (Kingma & Welling, 2014; Rezende et al., 2014), the optimization of TreeVAE ultimately comes down to maximizing the ELBO. As shown by Manduchi et al. (2023), the ELBO in this setting can be written as follows:
@@
Hereby, we distinguish between the reconstruction term and the KL divergence term between the variational posterior and the prior of the tree, which can be further decomposed into contributions from the root, the remaining nodes, and decisions, as indicated in Equation equation 23. Both terms are approximated using Monte Carlo (MC) sampling during training, as elaborated further in Manduchi et al. (2023).
@@
To grow the binary tree structure $\mathcal{T}$, TreeVAE begins with a simple tree configuration, typically composed of a root and two leaves. This initial structure is trained for a defined number of epochs, optimizing the ELBO. Subsequently, the model iteratively expands the tree by attaching two new child nodes to a current leaf node in the model. In their approach, Manduchi et al. (2023) opted to expand the tree by selecting nodes with the highest number of assigned samples, thus implicitly encouraging balanced leaves. The sub-tree formed by the new leaves and the parent node undergoes training for another number of epochs, keeping the weights of the remaining model frozen. This expansion process continues until either the tree reaches its maximum depth $H$, a predefined maximum number of effective leaves, or another predefined condition is met. Optionally, after the tree has been expanded, all parameters in the model may be fine-tuned for another predefined number of epochs. Finally, the tree is pruned to remove empty branches.
@@
To improve the clustering performance, especially for colored images, Manduchi et al. (2023) enhance TreeVAE with contrastive learning. This addition enables the model to encode prior knowledge on data invariances through augmentations, facilitating the learning process and better capturing meaningful relationships within complex data. By incorporating contrastive objectives into the training process, TreeVAE becomes more adept at retrieving semantically meaningful clusters from colored image data.
Appendix B Diffusion Models
@@
Diffusion models have garnered considerable attention in recent years for their remarkable image-generation capabilities, outperforming traditional GANs in achieving high-quality results. This surge in interest has led to the development of numerous diffusion-based models, with the initial concept being introduced by Sohl-Dickstein et al. (2015), drawing inspiration from principles rooted in thermodynamics. In this section, we delve into one of the first and most well-known diffusion models, namely the DDPM, outlined in Section B.1. Additionally, we explore one possible integration of DDPM with VAEs to leverage the interpretable latent space offered by VAEs along with the superior generation quality of DDPM. Note that we use the DDIM instantiation at inference time.
B.1 Diffusion Denoising Probabilistic Models
@@
Diffusion Denoising Probabilistic Models (DDPM), introduced by Ho et al. (2020), are latent variable models that consist of two opposed processes. The main idea behind DDPM involves iteratively adding noise to the original image, thereby progressively degrading the signal with each step until only noise remains. In a second step, the image is successively reconstructed through a denoising process, employing a learned function to remove the noise. Therefore, DDPM essentially entail a forward noising process followed by a reverse denoising process, aiming to restore the original image from its noisy counterpart. Subsequently, both these processes will be explored in more detail.
@@
The forward process, also known as the diffusion process or forward noising process, serves a similar purpose as the inference model in VAEs. It operates as a Markov chain that gradually introduces noise to the data signal $\bm{x}_{0}$ over $T$ steps, where the intermediate states of the deformed data are denoted as $\bm{x}_{t}$ for $t=1,\dots,T$. This process follows a trajectory determined by a noise schedule ${\beta_{1},\dots,\beta_{T}}$, which controls the rate at which the original data is degraded. While it is possible to learn the variances $\beta_{t}$ via reparametrization, they are often chosen as hyperparameters with a predetermined schedule (Ho et al., 2020). Assuming Gaussian noise, the forward process can be represented as follows:
@@
[…]
@@
In summary, the reverse process models a complex target data distribution by sequentially transforming simple distributions through a generative Markov chain. This approach overcomes the traditional tradeoff between tractability and flexibility in probabilistic models (Sohl-Dickstein et al., 2015), leading to a flexible and efficient generative model:
@@
Ho et al. (2020) keep the time dependent variances $\bm{\Sigma}_{\theta}\left(\bm{x}_{t},t\right)=\sigma_{t}^{2}\bm{I}$ constant, typically choosing $\sigma_{t}^{2}=\beta_{t}$ or $\sigma_{t}^{2}=\tilde{\beta}_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$. Therefore, only the time-dependent posterior mean function $\bm{\mu}_{\theta}$ is learned during training. In fact, Ho et al. (2020) further suggest the following reparametrization of the posterior means:
$\bm{\mu}_{\theta}\left(\bm{x}_{t},t\right)=\tilde{\bm{\mu}}_{t}\left(\bm{x}_{t},\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(\bm{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{\theta}\left(\bm{x}_{t}\right)\right)\right)=\frac{1}{\sqrt{\alpha_{t}}}\left(\bm{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\bm{\epsilon}_{\theta}\left(\bm{x}_{t},t\right)\right),$ (31)
@@
$\mathcal{L}=\mathbb{E}_{q}[\underbrace{\operatorname{KL}\left(q\left(\bm{x}_{T}\mid\bm{x}_{0}\right)\|p\left(\bm{x}_{T}\right)\right)}_{\mathcal{L}_{T}}+\sum_{t>1}\underbrace{\operatorname{KL}\left(q\left(\bm{x}_{t-1}\mid\bm{x}_{t},\bm{x}_{0}\right)\|p_{\theta}\left(\bm{x}_{t-1}\mid\bm{x}_{t}\right)\right)}_{\mathcal{L}_{t-1}}-\underbrace{\log p_{\theta}\left(\bm{x}_{0}\mid\bm{x}_{1}\right)}_{\mathcal{L}_{0}}]$ (32)
For small diffusion rates $\beta_{t}$, the forward and the reverse processes share the same functional form, following Gaussian distributions (Sohl-Dickstein et al., 2015). Thus, the KL divergence terms between the posteriors of the forward process equation 27 and the reverse process equation 29 have a closed form. Ho et al. (2020) further simplified the training process by introducing a more efficient training objective. Given the assumption of fixed variances, $\mathcal{L}_{t-1}$ can be written as:
@@
where $C$ is a constant that does not depend on the model parameters $\theta$. Thus, the reverse process mean function, $\bm{\mu}_{\theta}$, is optimized to predict $\tilde{\bm{\mu}}_{t}$, the fixed noisy mean function at time step $t$ from the forward process equation 28. However, instead of directly comparing $\bm{\mu}_{\theta}$ and $\tilde{\bm{\mu}}_{t}$, by reparametrization, the model can be trained to predict the noise $\bm{\epsilon}$ at any given time step $t$ which results in more stable training results according to Ho et al. (2020). Thus, equation 33 can be further rewritten as
@@
where $\bm{\epsilon}_{\theta}$ is a function parametrized as a neural network that maintains equal dimensionality for input and output. The preferred architecture for $\bm{\epsilon}_{\theta}$ is a U-Net (Ronneberger et al., 2015), as used by Ho et al. (2020), Dhariwal & Nichol (2021), and Pandey et al. (2022). The training of $\bm{\epsilon}_{\theta}$ involves multiple epochs, where for each sample $\bm{x}_{0}$ of the original data, a time step $t$ within the diffusion sequence $1,\dots,T$ is chosen at random. Subsequently, some noise $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\bm{I})$ is sampled for that specific time step, which needs to be predicted by $\bm{\epsilon}_{\theta}$. To optimize the model, a gradient step is computed with respect to the loss function detailed in Equation equation 34. Therefore, this training procedure does not require iterating through every step of the diffusion model, circumventing the issue of slow sample generation that is typical for diffusion models.
Appendix C Additional Qualitative Results
C.1 Generations on MNIST

### v2

@@ Appendix 0.A Detailed Formulation of TreeVAE
Here, we provide a more in-depth introduction to TreeVAE, closely following the original paper [19]. Built upon the VAE framework [14], TreeVAEs inherently possess stochastic latent variables. However, unlike traditional VAEs, TreeVAEs organize these latent variables in a learnable binary tree structure as shown in Figure 7, enabling them to capture complex hierarchical relationships.
@@
In the TreeVAE framework, the tree hierarchically divides the data, yielding separate stochastic embeddings for each node. This division into nodes induces a probability distribution, allowing each sample to navigate through the nodes of the tree in a probabilistic manner, from the root to the leaf. Ideally, high-variance features should be partitioned at earlier stages in the tree, while deeper nodes encapsulate more detailed concepts. The flexible tree structure is learned during training and is specific to the data distribution. This structure allows for the incorporation of sample-specific probability distributions over the different paths in the tree, as further explained in Section 0.A.1. In this manner, TreeVAEs contribute to the generation of comprehensive hierarchical data representations.
@@
TreeVAE combines elements from both LadderVAE [32] and ANT [33], with the aim to create a VAE-based model capable of performing hierarchical clustering via the latent variables. Similar to LadderVAE, the inference and the generative model share the same top-down hierarchical structure. In fact, if we isolate a path from the root to any leaf in the tree structure of TreeVAE, we obtain an instance of the top-down model seen in LadderVAE. On the other side, both Adaptive Neural Trees and TreeVAE engage in representation and architecture learning during training. However, while ANTs are tailored for regression and classification tasks, TreeVAE’s focus lies in hierarchical clustering and generative modeling. This allows TreeVAEs to generate new class-specific data samples based on the latent embeddings of the leaves from the tree.
0.A.1 Model Formulation
@@
TreeVAE comprises an inference model and a generative model, which share the global structure of the binary tree $\mathcal{T}$. Following the original work [19], the tree structure is learned during training using dataset $\bm{X}$ and a predetermined maximum tree depth denoted as $H$. Specifically, the tree structure entails the set of nodes $\mathbb{V}=\{0,...,V\}$, the subset of leaves $\mathbb{L}\subset\mathbb{V}$, and the set of edges $\mathcal{E}$.
@@
The transformations and routers are learned during the model training process, as further elaborated in Section 0.A.2, which comprehensively covers all aspects of model training. The exact parametrizations of the transformations and routers vary depending on whether they are applied in the context of the inference or generative model. Subsequently, the specifics of both of these models will be explored.
@@
Following the generation of latent representations for each node $i$ in the tree, the decision regarding traversal to the left or right child node are determined based on the sampled $\bm{z}_{i}$ value. These decisions, denoted by $\mathrm{c}_{i}$, for each non-leaf node $i$ are guided by routers, which are neural network functions linked to the generative model and defined as $\left\{r_{p,i}\mid i\in\mathbb{V}\backslash\mathbb{L}\right\}$. Here, $\mathrm{c}_{i}=0$ indicates the selection of the left child of the internal node $i$. As mentioned before, these decisions follow a Bernoulli distribution, $\mathrm{c}_{i}\mid\bm{z}_{i}\sim\operatorname{Ber}\left(r_{p,i}\left(\bm{z}_{i}\right)\right),$ which, in turn, induces the following path probability between a node $i$ and its parent node $pa(i)$:
@@
To conclude the generative process, $\bm{x}$ is obtained based on the latent embeddings of a selected leaf $l$. Nevertheless, assumptions on the distribution of the inputs are required. For real-valued $\bm{x}$, such as in colored datasets, [19] assume a Gaussian distribution. For grayscale images, they consider the Bernoulli distribution. Thus,
@@
where $\left\{\mu_{x,l},\sigma_{x,l}\mid l\in\mathbb{L}\right\}$ or $\left\{\mu_{x,l}\mid l\in\mathbb{L}\right\}$, respectively, are implemented as leaf-specific neural networks called “decoders". Typically, a simplification is made by assuming $\sigma_{x,l}^{2}\left(\bm{z}_{l}\right)=\bm{I}$ for convenience.
@@
The inference model of TreeVAE, depicted on the left side of Figure 7, introduces a deterministic bottom-up pass to incorporate the conditioning on $\bm{x}$, distinguishing it from the generative model. This bottom-up process is akin to the framework proposed by LadderVAE [32]. In this hierarchical structure, the bottom-up deterministic variables depend on each other via a series of neural network operations, specifically MLP of the same architecture as the transformation MLP defined previously,
@@
The first deterministic variable in this chain, denoted as $\mathbf{d}_{H}$, serves as the output of an encoder neural network. This encoder accepts the input image $\bm{x}$ and transforms it into the flattened, low-dimensional vector $\mathbf{d}_{H}$. $H$ corresponds to both the number of deterministic variables involved in the bottom-up process and the maximum depth to which the tree structure can grow during training as further explained in Section 0.A.2.
The tree structure is shared between the inference and the generative model, though adjustments are made to the node-specific parameterizations of the Gaussian distributions. Notably, in the inference phase, information is introduced through conditioning on $\bm{x}$, influencing the distribution of latent embeddings across all nodes. Hence, similar to LadderVAE, the means $\bm{\mu}_{q,i}$ and variances $\bm{\sigma}_{q,i}^{2}$ of the variational posterior distribution for each node $i$ are calculated using dense linear network layers conditioned on the deterministic variable of the same depth as the node $i$:
@@
Finally, given the variational distributions of the latent embeddings equation 16 & equation 17 and the variational distributions of the decisions equation 18, the variational posterior distribution of the latent embeddings and paths corresponds to:
[…]
0.A.2 Model Training
TreeVAE [19] entails the training of various components, involving both the parameters of the neural network layers present in the inference and generative models, as detailed in Section 0.A.1, and the binary tree structure $\mathcal{T}$. The training process alternates between optimizing the model parameters with a fixed tree structure and expanding the tree.
[…]
@@ 0.A.2 Model Training
Similar to other Variational Autoencoders [14], the optimization of TreeVAE ultimately comes down to maximizing the ELBO. As shown by [19], the ELBO in this setting can be written as follows:
@@
Hereby, we distinguish between the reconstruction term and the KL divergence term between the variational posterior and the prior of the tree, which can be further decomposed into contributions from the root, the remaining nodes, and decisions, as indicated in Equation equation 21. Both terms are approximated using Monte Carlo (MC) sampling during training, as elaborated further in [19].
@@
To grow the binary tree structure $\mathcal{T}$, TreeVAE begins with a simple tree configuration, typically composed of a root and two leaves. This initial structure is trained for a defined number of epochs, optimizing the ELBO. Subsequently, the model iteratively expands the tree by attaching two new child nodes to a current leaf node in the model. In their approach, the authors [19] opted to expand the tree by selecting nodes with the highest number of assigned samples, thus implicitly encouraging balanced leaves. The sub-tree formed by the new leaves and the parent node undergoes training for another number of epochs, keeping the weights of the remaining model frozen. This expansion process continues until either the tree reaches its maximum depth $H$, a predefined maximum number of effective leaves, or another predefined condition is met. Optionally, after the tree has been expanded, all parameters in the model may be fine-tuned for another predefined number of epochs. Finally, the tree is pruned to remove empty branches.
@@
Contrastive learning is used to improve the clustering performance, especially for colored images. This addition enables the model to encode prior knowledge on data invariances through augmentations, facilitating the learning process and better capturing meaningful relationships within complex data. By incorporating contrastive objectives into the training process, TreeVAE becomes more adept at retrieving semantically meaningful clusters from colored image data.
Appendix 0.B Detailed Formulation of Diffusion Models
@@
Diffusion models have garnered considerable attention in recent years for their remarkable image-generation capabilities, outperforming traditional GANs in achieving high-quality results. This surge in interest has led to the development of numerous diffusion-based models, with the initial concept being introduced by [30], drawing inspiration from principles rooted in thermodynamics. In this section, we delve into one of the first and most well-known diffusion models, namely the DDPM, outlined in Section 0.B.1. Additionally, we explore one possible integration of DDPM with VAEs to leverage the interpretable latent space offered by VAEs along with the superior generation quality of DDPM. Note that we use the DDIM instantiation at inference time.
0.B.1 Diffusion Denoising Probabilistic Models
@@
Diffusion Denoising Probabilistic Models (DDPM) [11] are latent variable models that consist of two opposed processes. The main idea behind DDPM involves iteratively adding noise to the original image, thereby progressively degrading the signal with each step until only noise remains. In a second step, the image is successively reconstructed through a denoising process, employing a learned function to remove the noise. Therefore, DDPM essentially entail a forward noising process followed by a reverse denoising process, aiming to restore the original image from its noisy counterpart. Subsequently, both these processes will be explored in more detail.
@@
The forward process, also known as the diffusion process or forward noising process, serves a similar purpose as the inference model in VAEs. It operates as a Markov chain that gradually introduces noise to the data signal $\bm{x}_{0}$ over $T$ steps, where the intermediate states of the deformed data are denoted as $\bm{x}_{t}$ for $t=1,\dots,T$. This process follows a trajectory determined by a noise schedule $\beta_{1},\dots,\beta_{T}$}, which controls the rate at which the original data is degraded. While it is possible to learn the variances $\beta_{t}$ via reparametrization, they are often chosen as hyperparameters with a predetermined schedule [11]. Assuming Gaussian noise, the forward process can be represented as follows:
@@
$\text{ where }\quad\tilde{\bm{\mu}}_{t}\left(\bm{x}_{t},\bm{x}_{0}\right)$ $=\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\bm{x}_{0}+\frac{\sqrt{\alpha_{t}}\left(1-\bar{\alpha}_{t-1}\right)}{1-\bar{\alpha}_{t}}\bm{x}_{t}\quad\text{ and }\quad\tilde{\beta}_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}.$ (26)
@@
In summary, the reverse process models a complex target data distribution by sequentially transforming simple distributions through a generative Markov chain. This approach overcomes the traditional tradeoff between tractability and flexibility in probabilistic models [30], leading to a flexible and efficient generative model:
@@
[11] keep the time dependent variances $\bm{\Sigma}_{\theta}\left(\bm{x}_{t},t\right)=\sigma_{t}^{2}\bm{I}$ constant, typically choosing $\sigma_{t}^{2}=\beta_{t}$ or $\sigma_{t}^{2}=\tilde{\beta}_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$. Therefore, only the time-dependent posterior mean function $\bm{\mu}_{\theta}$ is learned during training. In fact, [11] further suggest the following reparametrization of the posterior means:
$\bm{\mu}_{\theta}\left(\bm{x}_{t},t\right)$ $=\tilde{\bm{\mu}}_{t}\left(\bm{x}_{t},\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(\bm{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{\theta}\left(\bm{x}_{t}\right)\right)\right)$
$=\frac{1}{\sqrt{\alpha_{t}}}\left(\bm{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\bm{\epsilon}_{\theta}\left(\bm{x}_{t},t\right)\right),$
@@
$\mathcal{L}=$ $\mathbb{E}_{q}[\underbrace{\operatorname{KL}\left(q\left(\bm{x}_{T}\mid\bm{x}_{0}\right)\|p\left(\bm{x}_{T}\right)\right)}_{\mathcal{L}_{T}}$
$+\sum_{t>1}\underbrace{\operatorname{KL}\left(q\left(\bm{x}_{t-1}\mid\bm{x}_{t},\bm{x}_{0}\right)\|p_{\theta}\left(\bm{x}_{t-1}\mid\bm{x}_{t}\right)\right)}_{\mathcal{L}_{t-1}}$
$-\underbrace{\log p_{\theta}\left(\bm{x}_{0}\mid\bm{x}_{1}\right)}_{\mathcal{L}_{0}}]$
For small diffusion rates $\beta_{t}$, the forward and the reverse processes share the same functional form, following Gaussian distributions [30]. Thus, the KL divergence terms between the posteriors of the forward process equation 25 and the reverse process equation 27 have a closed form. The training process can be further simplified [11] by introducing a more efficient training objective. Given the assumption of fixed variances, $\mathcal{L}_{t-1}$ can be written as:
@@
where $C$ is a constant that does not depend on the model parameters $\theta$. Thus, the reverse process mean function, $\bm{\mu}_{\theta}$, is optimized to predict $\tilde{\bm{\mu}}_{t}$, the fixed noisy mean function at time step $t$ from the forward process equation 26. However, instead of directly comparing $\bm{\mu}_{\theta}$ and $\tilde{\bm{\mu}}_{t}$, by reparametrization, the model can be trained to predict the noise $\bm{\epsilon}$ at any given time step $t$ which results in more stable training results according to [11]. Thus, equation 29 can be further rewritten as
@@
where $\bm{\epsilon}_{\theta}$ is a function parametrized as a neural network that maintains equal dimensionality for input and output. The preferred architecture for $\bm{\epsilon}_{\theta}$ is a U-Net [26], as used by [11, 5, 23]. The training of $\bm{\epsilon}_{\theta}$ involves multiple epochs, where for each sample $\bm{x}_{0}$ of the original data, a time step $t$ within the diffusion sequence $1,\dots,T$ is chosen at random. Subsequently, some noise $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\bm{I})$ is sampled for that specific time step, which needs to be predicted by $\bm{\epsilon}_{\theta}$. To optimize the model, a gradient step is computed with respect to the loss function detailed in Equation equation 30. Therefore, this training procedure does not require iterating through every step of the diffusion model, circumventing the issue of slow sample generation that is typical for diffusion models.
Appendix 0.C Additional Qualitative Results
0.C.1 Generations on CIFAR-10